I have to translate problems like these into a system of equations, then solve it
ex:
The sum of two numbers is 36. The difference of one sixth of the larger number and one ninth of the smaller number is 11. Find the numbers.

Found 2 solutions by scott8148, ewatrrr:
Answer by scott8148(6628)

"The sum of two numbers is 36" ___ L + s = 36 ___ 4L + 4s = 144

"The difference of one sixth of the larger number and one ninth of the smaller number is 11"
___ (L / 6) - (s / 9) = 11

multiplying by 36 (LCD) ___ 6L - 4s = 396

adding equations (to ELIMINATE s) ___ 10L = 540

Answer by ewatrrr(24785)


Hi
The sum of two numbers is 36
Let x and (36-x)represent the smaller and larger number respectively
Question states***
(1/6)(36-x) + x/9 = 11
Solving for x
3(36-x) + 2x = 11*18
108 - 198 = x
-90 = x, smaller number. Larger number is 126. (36-(-90)
CHECKING our Answer***
(1/6)126 - 10 = 21 - 10 = 11